# Zhou, Orfanoudaki, and Zhu

### v2

@@ 11.1 Computation Resources
The computational resources used in our experiments are detailed in Table 1. Our code is implemented in Python, with key dependencies including Scikit-learn (pedregosa2011scikit) for solving linear programs and working with Gaussian mixture models (fitting and sampling), and CDD (fukuda1997cdd) for computing the vertices of polytopes. We also use cvxpy (diamond2016cvxpy) and cvxpylayers (cvxpylayers2019) libraries to solve convex optimization problems and implement differentiable optimization layers. We use the default parameter settings in their libraries across all experiments. A complete list of dependencies and their version numbers is available in our codebase.
11.2 Synthetic Setting
@@
We denote $\sigma$ as the component variance scale ($\sigma=1$ by default), and we denote $\mathbf{I}_{2}$ as the two-dimensional identity matrix.
Setting I
@@
A linear programming problem featuring a triangular feasible region with three vertex decisions, defined as:
$\max_{z\in\mathbb{R}^{2}}\left\{Y_{1}z_{1}+Y_{2}z_{2}\mid z_{1}+z_{2}\leq 1,z_{i}\geq 0,z_{2}\geq 0\right\},\quad\begin{pmatrix}Y_{1}\\ Y_{2}\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}-1\\ -1\end{pmatrix},\sigma\cdot\mathbf{I}_{2}\right).$
@@
This optimization problem can be interpreted as a profit maximization task, where a manufacturer chooses the optimal production quantities $z_{1}$ and $z_{2}$ under a budget constraint. The Gaussian random revenues $Y_{1}$ and $Y_{2}$ have negative expected values but may exhibit some variance, capturing a risky market scenario that could still yield profit under favorable conditions.
The feasible region in this problem can be more compactly denoted by its constraint matrix $\mathbf{A}$ and constraint vector $\mathbf{b}$ as
$\mathbf{A}=\begin{pmatrix}1&1\\ -1&0\\ 0&-1\end{pmatrix}\quad\mathbf{b}=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$.
Setting II
@@
A linear programming problem that employs a more complex octagonal feasible region with five vertices and multimodal objective uncertainty, allowing us to assess performance in scenarios with multiple potentially optimal decisions. Specifically, the optimization problem is set up as the canonical form $\max_{z\in\mathbb{R}^{2}}\left\{Y^{\top}z\mid\mathbf{A}z\leq\mathbf{b}\right\}$, where the constraint matrices $\mathbf{A}$ and constraint vector $\mathbf{b}$ are defined as
$\mathbf{A}=\begin{pmatrix}-0.5&-1\\ 0&-1\\ -0.5&1\\ 0.5&1\\ 2&-1\\ 1&0\\ 0&1\\ -1&0\end{pmatrix}\quad\mathbf{b}=\begin{pmatrix}-1\\ 0\\ 1\\ 5\\ 10\\ 5.5\\ 2.5\\ -1\end{pmatrix}$.
@@
where the mixture weights are $\boldsymbol{w}=(0.3,0.4,0.3),$ the component means are
@@
and the component variances are
11.3 Real-World Setting
As described in the main paper, we consider a real-world power grid investment decision-making problem (zhou2024hierarchical). We begin by introducing the application background, followed by a formalization of the problem as a knapsack optimization. Finally, we present a linear programming relaxation of this problem, which can be directly used in CREDO.
Background
A utility company based in Indianapolis, Indiana, has compiled detailed records of over 1,700 solar panel installations between 2010 and 2024, including the installation dates and affiliated grid components. With the renewable energy sector now at full scale, the management team anticipates a steady and significant monthly increase in solar adoption in the downtown area. In preparation for the incoming demand, they are planning targeted upgrades to grid-level inverters at four selected substations (we refer to them as Substation A to D) under a limited budget. The utility company would like to consult on CREDO’s suggestion for a list of candidate upgrade plans that would most likely be optimal.
The data is available in our codebase in a spatio-temporally aggregated format (monthly, by substation), with substation names anonymized. However, the granular solar panel installation data used in this study cannot be shared publicly, as it is proprietary to the utility company.
Mathematical Formulation
This problem can be mathematically formulated as a penalized knapsack problem, whose goal is to minimize the total penalty from capacity violations while ensuring that the total upgrade cost does not exceed the budget. Formally, define $d$ as the total number of substations, and define parameters as in Table 2, the optimization problem can be written as
$\min_{\mathbf{a}\in\{0,1\}^{d}}\sum_{i=1}^{d}\mathbbm{1}(Y_{i}\geq\tau_{i})(1-a_{i})l_{i}\quad\text{s.t.}\quad\sum_{i=1}^{d}a_{i}c_{i}\leq b.$ (26)
Table 2: Parameter definitions for the real-world optimization problem
Parameter Definition $a_{i}$ Binary decision variable; $a_{i}=1$ indicates a substation upgrade, and $a_{i}=0$ indicates no upgrade. $Y_{i}$ Random variable representing the monthly increase in solar panel installations. $c_{i}$ Cost associated with upgrading substation $i$. $\tau_{i}$ Capacity threshold—the maximum number of allowable solar connections at substation $i$. $l_{i}$ Penalty incurred if the threshold $\tau_{i}$ is exceeded without an upgrade. $b$ Total budget.
In our experiment, we configure these parameters as follows: we assume the cost $c_{i}$ and loss $l_{i}$ for all substations equals one unit. The capacity threshold $\tau_{i}$ is set to be the historical average solar panel monthly increment. The budget $b$ is set to half the cost of upgrading all substations, allowing at most two out of four substations can be upgraded. Note that these parameters require additional information from the company, which is not available at the time of writing this paper.
Relaxed Formulation
The knapsack problem defined in (26) can be relaxed into the following linear programming problem:
$\min_{z\in[0,1]^{d}}\sum_{i=1}^{d}\frac{l_{i}}{1+e^{-\beta(Y_{i}-\tau_{i})}}(1-a_{i})\quad\text{s.t.}\quad-\sum_{i=1}^{d}c_{i}z_{i}\leq b-\sum_{i=1}^{d}c_{i},$ (27)
where $\beta$ is an arbitrary smoothing parameter that we set $\beta=0.5$. Note that in the equation above, two relaxations have been applied:
  • The binary random variable $\mathbbm{1}\left\{Y_{i}\geq\tau_{i}\right\}$ has been relaxed to a continuous random variable over $[0,1]$ using the sigmoid function, where the smoothing parameter $\beta$ controls the sharpness of the approximation.
  • The decision space of $a_{i}$ has been relaxed from the discrete set $\{0,1\}^{d}$ to the continuous unit hypercube $0,1]^{d}$. This relaxation does not introduce approximation bias, as the optimal solutions to the original and relaxed problems are almost surely the same – both attained at one of the hypercube’s vertices.
Denote the vectors
$Y=\begin{pmatrix}\frac{l_{i}}{1+e^{-\beta(Y_{i}-\tau_{i})}}\\ \ldots\\ \frac{l_{i}}{1+e^{-\beta(Y_{i}-\tau_{i})}}\end{pmatrix},\quad\mathbf{c}=\begin{pmatrix}c_{1}\\ \ldots\\ c_{n}\end{pmatrix},\quad z=\begin{pmatrix}1-a_{1}\\ \ldots\\ 1-a_{n}\end{pmatrix}$, (28)
we can compactly write the optimization problem above as
$\min_{z\in\mathbb{R}^{d}}Y^{\top}z\quad\text{s.t.}\quad-\mathbf{c}^{\top}z\leq b-\mathbf{1}_{d}^{\top}\mathbf{c},~\mathbf{0}\leq z\leq\mathbf{1}.$ (29)
This can be written as the canonical form of linear programming $\min_{z\in\mathbb{R}^{d}}\left\{Y^{\top}z\mid\mathbf{A}z\leq\mathbf{b}\right\}$ by defining
$\mathbf{A}=\begin{pmatrix}-\mathbf{c}^{\top}\\ \mathbf{I}_{d}\\ -\mathbf{I}_{d}\end{pmatrix},\quad\mathbf{b}=\begin{pmatrix}b-\mathbf{1}_{d}^{\top}\mathbf{c}\\ \mathbf{1}_{d}\\ \mathbf{0}_{d}\end{pmatrix}$,
where $\mathbf{I}_{d}$ denotes the $d$-dimensional identity matrix, and $\mathbf{1}_{d}$ and $\mathbf{0}_{d}$ denote the $d$-dimensional all-ones and all-zeros vectors, respectively. These parameters can be directly input to the CREDO framework using the data $Y$ obtained via sigmoid transformation defined in (28) to obtain decision risk estimates.
11.4 Benchmarking Baselines Description
In this part, we describe the baselines, metrics, as well as the detailed evaluation procedure for each experiment that we conducted in the main paper. For both experiments, the following (Table 3) default values of hyperparameters are used for CREDO.
Table 3: Hyperparameters in the Convservativeness vs. Accuracy Tradeoff experiment.
Hyperparameter Description Default value $K$ Simulation size $1\times 10^{2}$ $\sigma$ (Component) variance scale 11 $n$ Calibration dataset size $1\times 10^{1}$ $m$ Training dataset size $1\times 10^{1}$
Table 4 summarizes the evaluation metrics used in our experiment. We use $T$ to denote the total number of repeated independent trials, where we set $T=20$ across all of our experiments. For each metric, the error is reported as the standard error of the mean (SEM), calculated as the standard deviation of the metric values across trials divided by $\sqrt{T}$. Note that while most metrics require summing over all feasible decisions, we can simplify to summing all the decisions defined on the vertices of the feasible region, as we have justified in the main text.
Table 4: Summary table of metrics used in the Convservativeness vs. Accuracy Tradeoff experiment.
Metric Name Description Mathematical Formula Validity The percentage of decisions where the risk estimate satisfies the conservativeness guarantee in (2) averaged across all trails. 1T ⋅—V— ∑_z ∈V 1 { ^α(z) ≤α(z)} TPR The percentage of decisions where the method correctly identifies them to have risk of less than one, averaged across all trials. 1T ⋅—V— ∑_z ∈V 1 { ^α(z) ¡ 1 } ⋅1 { α(z) ¡ 1 } MAE The mean absolute error of the estimated compared to the true risks across all decisions and trials. 1T ⋅—V— ∑_z ∈V — ^α(z) - α(z) —
@@ 11.4 Benchmarking Baselines Description
Table 5 presents the four selected baselines used in the Decision Quality Evaluation experiment. The baselines have been selected to satisfy the following criteria: (i) the ability to produce a decision that maximizes a (linear) objective function; and (ii) the capacity to handle randomness in the underlying optimization problem.
Table 5: Summary table of baselines used in the Decision Quality Evaluation experiment
Model Name Description Additional remarks PTO The standard two-stage predict-then-optimize approach (bertsimas2020predictive), which first predicts parameters and then solves the resulting optimization problem; min_z ∈Z f(z; ^Y), where $\hat{Y}$ is a point estimate of $\mathbb{E}[Y|X]$. Since $X$ is omitted in our setting, PTO degenerates to stochastic optimization, and $\hat{Y}$ can be simply specified as the empirical average estimator. RO Robust optimization defined as min_z ∈Z max_y ∈U f(z; y) where $\mathcal{U}\subseteq\mathcal{Y}$ is the uncertainty set of $Y$. The uncertainty set is constructed in a data-driven manner by applying naive CP to $Y$, using the $\ell_{\infty}$ norm residual as the nonconformity score (zhou2024hierarchical). SPO+ A predict-then-optimize method where the prediction model is trained using the surrogate smart predict-then-optimize loss (elmachtoub2022smart) For each trial, it is trained for $1\times 10^{2}$ epochs using a learning rate of $1\times 10^{-1}$. DFL Decision-focused learning, where the prediction model is trained by directly optimizing the downstream objective through end-to-end differentiation of the optimization layer (amos2017optnet). For each trial, it is trained for $1\times 10^{2}$ epochs using a learning rate of $1\times 10^{-1}$.
As for the evaluation metric, Algorithm 3 outlines the generic procedure for computing empirical confidence rankings in a single trial. In the synthetic setting, this procedure is repeated over $T=1\times 10^{2}$ independent trials with $\ell=1\times 10^{3}$ test data points. Meanwhile, in the real-world setting, the evaluation is performed rolling over $T=12$ periods, spanning from 2010 to 2022. Each trial uses a two-year window (24 months) of data, which is sequentially split into training, calibration, and testing sets in an 8:8:88~:~8~:~8 ratio. The reported error is the standard error of the mean (SEM), computed as the standard deviation of the metric across trials divided by $\sqrt{T}$.

### v3

@@ 10.1 Implementation Environments
All experiments were conducted on a machine running Windows 11, equipped with a 13th-generation Intel Core i7 CPU with 16 cores and 16 GB of RAM. No GPU acceleration was used in any of the experiments. The code is implemented in Python, and we list some key external packages: Scikit-learn (pedregosa2011scikit) is used for implementing Gaussian mixture models and some other statistical models that were used as baselines; CDD (fukuda1997cdd) is used for computing the vertices of polytopes to compute the closed-form solutions under the linear assumption; cvxpy and cvxpylayers (diamond2016cvxpy, cvxpylayers2019) are used to solve convex optimization problems and implement differentiable optimization layers; PyTorch is used to build gradient-based baseline models. Unless otherwise specified, we use the default parameter for all models in the package. The complete code and dependencies are available in our codebase.
10.2 Optimization problems
@@
We describe the details of the six optimization problems that are featured in our experiments. Throughout this section, we denote $\sigma$ as the component variance scale ($\sigma=1$ by default), and $\mathbf{I}_{d}$ as the $d$-dimensional identity matrix.
LP (Setting I)
@@
A linear programming problem with a triangular feasible region, defined as:
$\min_{z\in\mathbb{R}^{2}}Y_{1}z_{1}+Y_{2}z_{2}\quad\text{s.t.}\quad z_{1}+z_{2}\leq 1,z_{1}\geq 0,z_{2}\geq 0,$
@@
where $(Y_{1},Y_{2})^{\top}$ is a Gaussian random vector with mean $(-1,-1)^{\top}$ and covaraince matrix $\sigma\cdot\mathbf{I}_{2}$. This optimization problem can be interpreted as a profit maximization task, where a manufacturer chooses the optimal production quantities $z_{1}$ and $z_{2}$ under a budget constraint, and $Y_{1}$ and $Y_{2}$ represent a risky market scenario that could still yield profit under favorable conditions where the expected revenue is negative. Its feasible region can be more compactly represented in matrix form $\mathbf{A}z\leq\mathbf{b}$, where
$\mathbf{A}=\begin{pmatrix}1&1\\ -1&0\\ 0&-1\end{pmatrix},\quad\mathbf{b}=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$.
LP (Setting II)
@@
A linear programming problem that employs a more complex octagonal feasible region with five vertices and multimodal objective uncertainty, defined as
$\min_{z\in\mathbb{R}^{2}}Y_{1}z_{1}+Y_{2}z_{2}\quad\text{s.t.}\quad\mathbf{A}z\leq\mathbf{b},$
where the constraint matrix $\mathbf{A}$ and constraint vector $\mathbf{b}$ are defined as
$\mathbf{A}=\begin{pmatrix}-0.5&0&-0.5&0.5&2&1&0&-1\\ -1&-1&1&1&-1&0&1&0\end{pmatrix}^{\top},\quad\mathbf{b}=\begin{pmatrix}-1&0&1&5&10&5.5&2.5&-1\end{pmatrix}^{\top}$.
@@
where the mixture weights are $\boldsymbol{w}=(0.3,0.4,0.3),$ the component means $\boldsymbol{\mu}$ are defined as
$\mu_{1}=\begin{pmatrix}0.0,-0.8\end{pmatrix}^{\top},\quad\mu_{2}=\begin{pmatrix}-0.5,0.25\end{pmatrix}^{\top},\quad\mu_{3}=\begin{pmatrix}0.8,-0.1\end{pmatrix}^{\top}$,
and the component variances $\boldsymbol{\sigma}$ are defined as
Compared to the first setting, this setting allows us to assess risk assessment performance in more complex scenarios with multiple potentially optimal decisions.
QP
A quadratic programming problem defined as
$\min_{z\in\mathbb{R}^{2}}\frac{1}{2}z^{\top}Qz+\langle z,y\rangle\quad\text{s.t.}\quad\mathbf{A}z\leq\mathbf{b},$
where the parameters are defined similar to LP (Setting I):
$Q=\begin{pmatrix}0.1&0\\ 0&0.1\\ \end{pmatrix},\quad\mathbf{A}=\begin{pmatrix}1&1\\ -1&0\\ 0&-1\end{pmatrix},\quad\mathbf{b}=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$.
SOCP
A second-order conic programming problem, defined as
$\min_{z\in\mathbb{R}^{2}}~\langle z,y\rangle\quad\text{s.t.}\quad 0.1\cdot\|z\|_{2}\leq\mathbf{A}z-\mathbf{b}.$
where the parameters are defined similar to LP (Setting I):
$\mathbf{A}=\begin{pmatrix}1&1\\ -1&0\\ 0&-1\end{pmatrix},\quad\mathbf{b}=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$.
IP
An integer programming problem, defined as
$\min_{z\in\mathbb{Z}^{2}}~\langle z,y\rangle\quad\text{s.t.}\quad\mathbf{A}z-\mathbf{b}\leq 0.$
where the parameters are defined similar to LP (Setting II):
$\mathbf{A}=\begin{pmatrix}-0.5&0&-0.5&0.5&2&1&0&-1\\ -1&-1&1&1&-1&0&1&0\end{pmatrix}^{\top},\quad\mathbf{b}=\begin{pmatrix}-1&0&1&5&10&5.5&2.5&-1\end{pmatrix}^{\top}$.
Real Data
This setting considers a real-world power grid investment decision-making problem motivated by (zhou2024hierarchical). A utility company based in Indianapolis, Indiana, has compiled detailed records of over 1,700 solar panel installations between 2010 and 2024, including the installation dates and affiliated grid components. With the renewable energy sector now at full scale, the management team anticipates a steady and significant monthly increase in solar adoption in the downtown area. In preparation for the incoming demand, they are planning targeted upgrades to grid-level inverters at $d=4$ selected substations (we refer to them as Substation A to D) to minimize loss for grid failures subject to a limited budget. This decision-making problem can be formulated as a penalized knapsack problem, defined as
$\min_{\mathbf{a}\in\{0,1\}^{d}}\sum_{i=1}^{d}\mathbbm{1}(Y_{i}\geq\tau_{i})(1-a_{i})l_{i}\quad\text{s.t.}\quad\sum_{i=1}^{d}a_{i}c_{i}\leq b,$
where $a_{i}\in{0,1}$ is a binary decision variable indicating whether substation $i$ is upgraded ($a_{i}=1$) or not ($a_{i}=0$); $Y_{i}$ is a random variable representing the monthly increase in solar panel installations connected to substation $i$; $c_{i}$ denotes the cost associated with upgrading substation $i$; $\tau_{i}$ is the capacity threshold, corresponding to the maximum allowable number of solar connections at substation $i$; $l_{i}$ represents the penalty incurred when the realized installations exceed $\tau_{i}$ without an upgrade; and $b$ denotes the total available budget.
The optimization problem above can be further relaxed into an LP problem:
$\max_{z\in\mathbb{R}^{d}}~\langle Y,z\rangle\quad\text{s.t.}\quad\mathbf{A}z\leq\mathbf{b},$
where the parameters are defined as
$Y=\begin{pmatrix}{l_{1}}{\cdot(1+e^{-\beta(Y_{1}-\tau_{1})})^{-1}}\\ \ldots\\ {l_{d}}{\cdot(1+e^{-\beta(Y_{d}-\tau_{d})})^{-1}}\\ \end{pmatrix},\quad z=\begin{pmatrix}1-a_{1}\\ \ldots\\ 1-a_{d}\end{pmatrix},\quad\mathbf{A}=\begin{pmatrix}-\mathbf{c}^{\top}\\ \mathbf{I}_{d}\\ -\mathbf{I}_{d}\end{pmatrix},\quad\mathbf{b}=\begin{pmatrix}b-\mathbf{1}_{d}^{\top}\mathbf{c}\\ \mathbf{1}_{d}\\ \mathbf{0}_{d}\end{pmatrix}$.
They are manually configured as follows: The cost $c_{i}$ and loss $l_{i}$ for all substations equals one unit; The capacity threshold $\tau_{i}$ is set to be the historical average solar panel monthly increment; The budget $b$ is set to half the cost of upgrading all substations, allowing at most two out of four substations to be upgraded; The smoothing parameter $\beta$ is set to $0.5$, which controls the sharpness of approximation to the thresholding function $\mathbbm{1}(Y_{i}\geq\tau_{i})$. This relaxation is used in our experiment so that the derived closed-form solution could be applied.
Table 1: Hyperparameter settings for CREDO.
Hyperparameter Description Default value $K$ Simulation size $1\times 10^{2}$ $\sigma$ Component variance scale 11 $n$ Calibration dataset size $1\times 10^{1}$ $m$ Training dataset size $1\times 10^{1}$
Table 2: Summary table of selected metrics used in the experiments.
Metric Name Description Mathematical Formula Validity The percentage of decisions where the risk estimate satisfies the conservativeness guarantee in (2). 1—V— ∑_z ∈V 1 { ^α(z) ≤α(z)} MAE The mean absolute error of the estimated compared to the true risks. 1—V— ∑_z ∈V — ^α(z) - α(z) —
10.3 Baselines
In this part, we describe the baselines used in the experiments. To control for variables, the baselines share the same set of hyperparameters as CREDO, summarized in Table 1. Learning rates and their maximum epochs are set to $1\times 10^{-2}$ and $1\times 10^{2}$. All optimization procedures are run for $10$ epochs.
10.3.1 Baselines in Section 7.1
  • (ii)
    SA (sample average): A naive sample average using observed historic data,
@@ 10.3 Baselines
    Note that this estimator does not have any predictive power as it does not take $x$ as input, and does not guarantee conservativeness.
  • (iiii)
    LR (logistic regression): Models the probability of $z$ falling within $\pi(Y)$ via the logistic regression model:
[…]
    where $x$ denotes the feature vector and $(\theta,\theta_{0})$ are learned parameters.
  • (iiiiii)
[…]
  • (iviv)
    QE (quantile estimator): The $75\%$ empirical quantile of $\{\mathbbm{1}\{z\not\in\pi(y_{i})\}\}_{i=1}^{n}$. This is equivalent to quantile regression methods when $X$ is omitted.
  • ($v$)
    CP (conformal prediction): The naive conformal prediction method (vovk2005algorithmic) applied to $\{(x_{i},\mathbbm{1}\{z\in\pi(y_{i})\})\}_{i=1}^{n}$, taking the upper bound of the $25\%$ prediction interval as the final risk estimate:
[…]
    where $\hat{Q}_{e}(\cdot)$ denotes the empirical quantile of the nonconformity scores (i.e., regression prediction residuals). The regression model is taken as SA (the sample average estimator).
10.3.2 Baselines in Section 7.2
  • (ii)
[…]
  • (iiii)
    RO (robust optimization): a minmax optimization problem $\min_{z\in\mathcal{Z}}\max_{y\in\mathcal{U}}f(z;y),$ where the uncertainty set $\mathcal{U}$ is constructed using naive conformal prediction under the $\ell_{\infty}$ norm;
  • (iiiiii)
    SPO+ (smart predict-then-optimize): a predict-then-optimize method where the prediction model is trained using a surrogate loss that is convex and explicit (elmachtoub2022smart);
  • (iviv)
    DFL (decision-focused learning): prediction models are trained by directly optimizing the downstream objective through end-to-end differentiation of the optimization layer (cvxpylayers2019, amos2017optnet).
@@
The baselines are selected based on the following criteria: ($i$) The ability to produce a decision that maximizes a (linear) objective function; and ($ii$) The capacity to handle randomness in the underlying optimization problem.
10.3.3 Baselines in Section 7.3
We describe the baselines used for comparing the optimization procedures:
Algorithm 3 GD (Gradient Descent) for Solving (13b) in Generla Convex Settings
0: Prediction $\hat{y}$; Decision $z$; Penalty parameter $\lambda$; Learning rate $\eta$; Tolerance level $\epsilon$.
[…]
2:for $t\in\{1,\ldots,T\}$ do
3:  $\psi_{1}^{(t-1)}\leftarrow\nabla_{y}\min_{z^{\prime}\in\mathcal{Z}}f(z^{\prime};y)\big|_{y=y^{(t-1)}}$ compute via differentiable optimization;
4:  $\psi_{2}^{(t-1)}\leftarrow\nabla_{y}f(z;y)\big|_{y=y^{(t-1)}}$;
5:  $\psi^{(t-1)}\leftarrow\frac{y^{(t-1)}-\hat{y}}{\|y^{(t-1)}-\hat{y}\|_{2}}+\lambda\cdot\left(\psi_{1}-\psi_{2}+\epsilon\right)$ if constraint is positive else $\frac{y^{(t-1)}-\hat{y}}{\|y^{(t-1)}-\hat{y}\|_{2}}$;
6:  $y^{(t)}\leftarrow y^{(t-1)}-\eta\cdot\psi^{(t-1)}$;
7:end for
[…]
  • (ii)
    GD (gradient descent): A gradient descent-based algorithm that approximates (18) by minimizing its Lagrangian function, where its gradient is:
    $\nabla_{y}\mathcal{L}(y,\lambda)=\frac{y-\hat{y}}{\|y-\hat{y}\|_{2}}+\lambda\cdot\nabla_{y}\left[\min_{z^{\prime}\in\mathcal{Z}}f(z^{\prime};y)-f(z;y)+\epsilon\right]^{+}.$
    Here, $\lambda$ is the regularization hyperparameter, which is set to $1\times 10^{3}$. The term $\nabla_{y}\min_{z^{\prime}\in\mathcal{Z}}f(z^{\prime};y)$ is computed through differentiable optimization layers (cvxpylayers2019, amos2017optnet). The pseudocode of GD is provided in Algorithm 3.
  • (iiii)
    BF (brute force): A heuristic algorithm that first approximates the constraint space and decision space with gridded sets:
    $\tilde{\mathcal{Y}}\approx\{y\in\mathcal{Y}\mid f(z;y)-f(z^{\prime};y)>\epsilon\}\quad\text{and}\quad\tilde{\mathcal{Z}}\approx\mathcal{Z},$
    then solves the finite optimization problem $\min_{y\in\tilde{\mathcal{Y}}}\min_{z^{\prime}\in\tilde{\mathcal{Z}}}\|y-\hat{y}\|_{2}$.
We note that RS (random search) is similar to BF except that the gridded sets are constructed by drawing $T$ samples from a standard Gaussian distribution over the corresponding spaces.
10.4 Metrics
Table 2 summarizes three evaluation metrics used in our experiments. The reporting standards for the metrics are their means $\pm$ standard deviations across $T=100$ repeated trials.